# Distinguishing Tidal Disruption Events and Changing-look Active Galactic Nuclei via Variation of Mid-infrared Color

### v2

@@ Abstract
At present, there is a lack of effective probes to distinguish between mid-infrared (MIR) outbursts induced by tidal disruption events (TDEs) and changing-look active galactic nuclei (CLAGNs) based on only MIR data. Here, we propose that the time variation of MIR color is a promising probe. With an optically selected sample containing TDEs, ambiguous nuclear transients (ANTs), and CLAGNs, we studied the variation of MIR color after subtracting the quiescent fluxes using NEOWISE-R data. We found that the MIR color of TDEs and ANTs turns red faster than CLAGNs during the rising phase, and TDEs have a redder color than ANTs at the earliest phase. The former may be caused by the difference between the ultraviolet light curves of TDEs/ANTs and CLAGNs, or be related to no or weak underlying AGN in TDEs/ANTs, while the latter may be related to the difference in the dust geometry. Based on color variation rate, we selected high-probability TDE, ANT, and CLAGN candidates from MIR outbursts in samples of Jiang et al. (2021) and Masterson et al. (2024). We found that both samples are mixtures of TDEs/ANTs and CLAGNs. For MIR outbursts whose hosts are not Seyfert galaxies, we estimated that $\sim 50\%-80\%$ are TDEs and inferred a rate of infrared TDEs of $1.5-2.8\times 10^{-5}$ galaxy-1 yr-1, comparable with that of optical TDEs. The rest are CLAGNs, suggesting the presence of weak AGNs that cannot be identified using common diagnoses. Our work opens a new door for future classification of infrared selected transients.
@@ 1 Introduction
In quiescent galaxies, which account for the majority of galaxies, tidal disruption events (TDEs) are powerful probes of supermassive black holes (SMBHs). A TDE occurs when a star wanders too close to the SMBH for its own gravity to resist the tidal force (e.g., Rees, 1988). TDEs can be observed through bright flares lasting for months to years (e.g., van Velzen et al., 2020). Thanks to the development of time-domain surveys, $\gtrsim 100$ TDEs have been discovered. Most of TDEs are selected in the optical band (optical TDEs) via their bright UV-optical emission (e.g., van Velzen et al., 2020; Hammerstein et al., 2023), while a small number are selected in the X-ray band (X-ray TDEs) via their bright X-ray emission (e.g., Saxton et al., 2020; Sazonov et al., 2021), and some are bright in both bands (e.g., Guolo et al., 2024). According to recent sample studies, the incidence rate of optical TDEs is around $3\times 10^{-5}$ galaxy-1 yr-1 (e.g., Yao et al., 2023), while that of X-ray TDEs is around $1\times 10^{-5}$ galaxy-1 yr-1 (e.g., Sazonov et al., 2021). In addition, the TDE rate is related to the properties of the host galaxy. For example, post-starburst galaxies have a higher incidence rate of TDEs than normal galaxies (e.g., French et al., 2016; Yao et al., 2023).
@@
Despite these findings, there is a lack of complete TDE samples because dust-obscured TDEs are hard to detect by optical or X-ray surveys. Fortunately, the TDE’s UV emission heats the surrounding dust to produce infrared (IR) re-radiation, known as IR echo (e.g., Jiang et al., 2016; van Velzen et al., 2016). Dust-obscured TDEs can reveal themselves by their IR echoes as IR outbursts, for example, Arp 299-B AT1 (Mattila et al., 2018) and AT 2017gbl (Kool et al., 2020). Thanks to the all-sky mid-infrared (MIR) sky survey of the Wide-field Infrared Survey Explorer (WISE, Wright et al., 2010), hundreds of MIR outbursts in centers of galaxies have been discovered (Jiang et al., 2021; Reynolds et al., 2022; Masterson et al., 2024; Necker et al., 2024). The derived event rate is $\sim 2-5\times 10^{-5}$ galaxy-1 yr-1, similar to the rate of optical and X-ray TDEs, implying that optical and X-ray surveys may have missed a significant fraction of TDEs.
@@
However, a key question remains unanswered: are these MIR outbursts TDEs? Although the authors of previously mentioned works preferred that their samples consist mainly of obscured TDEs, there is doubt (e.g., Dodd et al., 2023) that changing-look active galactic nuclei (CLAGN) contribute to these MIR outbursts significantly. CLAGNs (see review in Ricci & Trakhtenbrot, 2023) can also produce MIR outbursts (e.g., Sheng et al., 2017; Lyu et al., 2022). Despite many differences between TDEs and CLAGNs in UV/optical bands in terms of light curves (LCs) and spectra (e.g., Zabludoff et al., 2021), distinguishing their IR echoes remains challenging. Differences in the UV/optical LCs are blurred by the light travel time effect of the echo, and the spectral features that may distinguishing the two are lost when the emission is reprocessed to the IR band. To make matters worse, there is a class of outbursts in active galactic nuclei (AGN) that share the characteristics of both TDEs and CLAGNs, for example AT 2016ezh (=PS16dtm, Blanchard et al., 2017; Petrushevska et al., 2023), AT 2017bgt (Trakhtenbrot et al., 2019), AT 2018bcb (=ASASSN-18jd, Neustadt et al., 2020), and some others. These outbursts are referred to as ambiguous nuclear transients (ANTs) as their natures are difficult to determine through UV/optical information: they may be TDEs that occur in AGNs (e.g., Blanchard et al., 2017), or enhanced accretions onto SMBHs (e.g., Trakhtenbrot et al., 2019). In any case, it is almost impossible to distinguish using MIR information only. The nature of MIR outbursts can be distinguished by radio jet (e.g. for Arp 299-B AT1, Mattila et al., 2018), broad emission lines in the near-infrared (NIR) spectra (e.g. for AT 2017gbl, Kool et al., 2020) or optical spectra (e.g., Wang et al., 2022a), or shape of X-ray spectrum (e.g., Masterson et al., 2024). However, these data are not available for the vast majority of MIR outbursts. Therefore, there needs to be an effective probe to distinguish between TDEs and CLAGNs based on only MIR data.
@@
Jiang et al. (2017) noticed that the dust temperature in AT 2016ezh was high with a blackbody temperature $T_{\rm BB}\gtrsim 2,000$ K initially, and then rapidly dropped to $T_{\rm BB}\sim 1,000$ K in 100–300 days. Although the temperature values here may not be accurate because the dust properties are uncertain, the trend of temperature decline is robust. Similar rapid drops in dust temperatures have also been observed in other TDEs, ANTs and TDE candidates (e.g., Mattila et al., 2018; Sun et al., 2020; Kool et al., 2020; Reynolds et al., 2022; Hinkle, 2024). However, such drops were not found in CLAGNs [^1] Although some works claimed a “redder when brighter” trend in MIR emission of CLAGNs (Yang et al., 2018), they did not subtract the host galaxy’s contribution to IR emission when they calculated the color. We guessed that there may be a general difference in the time variation of dust temperature between TDEs and CLAGNs. The primary purpose of this work is to verify this difference and attempt to distinguish TDEs from CLAGNs using this difference.
@@ 2 Optically-selected outbursts and data reduction
In order to study the variation of MIR colors of TDEs, CLAGNs and ANTs, and to check the disparity between different types, we collected outbursts whose type can be defined by optical observations. Although in the literature, the term ANT was also used by some authors to refer to different meanings, we will strictly use the term to mean the transient source occurring in an AGN whose nature is being debated. We required the outburst to have a bright MIR emission, and the variation of MIR color in its rising phase can be studied. The requirements and final samples are described in detail in sections 2.1 to 2.3, and the MIR properties of the sample are displayed in section 2.4.
@@ 2.1 MIR-bright TDEs, CLAGNs and ANTs
Lyu et al. (2022) has investigated the MIR properties of CLAGNs reported in the literature and found 68 CLAGNs with significant MIR variability. We started from Lyu’s sample but removed AT 2018dyk [^3] Referred to as SDSS J153308.02+443208.4 in Lyu et al. (2022) and SDSS J1115+0544. We generated the MIR LCs of the remaining 66 objects and selected bright objects with large amplitude variability in the same way as we described previously. One CLAGN, NGC 4151, is too bright, and its photometry was severely affected by saturated pixels, so we removed it from the sample. Finally, a total of 28 CLAGNs passed our criteria. As pointed out by Sheng et al. (2017), the CLAGNs we selected with large amplitudes of MIR variability are likely to originate from changes in the accretion rate, rather than changes in obscuration or other causes.
We collected ANTs from the literature and did the same thing. Eight ANTs passed our criteria, including AT 2016ezh, AT 2017bgt, AT 2018bcb, AT 2019aalc, AT 2019avd, AT 2020agdm, AT 2021loi and AT 2022fpx.
@@ 2.2 Quiescent and outburst states
The above analysis is based on LCs generated by NEOWISE single-exposure photometries. This method of generating LCs was proved to be reliable (e.g., Jiang et al., 2021); however, it leads to large magnitude errors for faint MIR sources. Thus, for objects with a quiescent W2 magnitude $>$11, we generated LCs using time-resolved coadded data provided by the unWISE project (Meisner et al., 2018) with an image subtraction method, which was also proved reliable (e.g., Masterson et al., 2024). For each object in each band, we made a reference image by combining the quiescent-state images, and then subtracted it from all the images to obtain difference images. We measured flux and error on each difference image using a PSF model whose center is fixed at the optical position for TDEs and ANTs or at the galaxy center for CLAGNs. In addition to routine statistical flux error, we considered the systematic error due to image subtraction, which is 2% of the quiescent PSF flux according to our tests using galaxies with no outbursts. We adopted the fluxes measured on difference images as excess fluxes $f_{e}$, and selected the outburst state in the same way as described previously. In Figure 2, we show an example that the data quality improves by using unWISE data. The flux errors are smaller, and more data points are selected for the outburst state, so the variation of MIR color can be better studied. However, because we have not obtained unWISE data in 2023 and beyond, we still use NEOWISE data for transients that started late, for example, AT 2022fpx, AT 2022upj, and AT 2023ugy.
@@ 2.3 Rising phase and further selections
WISEA J1003+3525 0.1189 CLAGN 26 $-0.06\pm 0.06$ $0.52\pm 0.10$
WISEA J1545+1709 0.0483 CLAGN 26 $-0.05\pm 0.10$ $0.71\pm 0.15$
@@ 2.4 MIR properties of the samples
[…]
@@ 3 Analysis of MIR color variation and results
This section studied the MIR color variation of the optically selected outbursts in the rising phase. We did not calculate the dust temperature as some previous works had done (e.g., Jiang et al., 2017; Reynolds et al., 2022). This is because the dust temperature depends heavily on the spectral energy distribution (SED), and its relationship with the directly observed quantity is complicated, making its error difficult to analyze. Instead, we used K-corrected MIR color $W1-W2$, i.e. color in the rest frame, as a probe of dust temperature. We describe how we applied the K-correction and estimated the systematic error introduced by the K-correction in Appendix B.

### v3

@@ Abstract
In this work, we propose that the time variation of mid-infrared (MIR) color is a promising probe to distinguish between MIR outbursts induced by tidal disruption events (TDEs) and changing-look active galactic nuclei (CLAGNs). With an optically selected sample containing TDEs, ambiguous nuclear transients (ANTs), and CLAGNs, we studied the variation of MIR color (W1W2W1-W2) after subtracting the quiescent fluxes using NEOWISE-R data. The MIR color of TDEs and ANTs turns red faster than CLAGNs during the rising phase, as the color variation rate (CVR) of TDEs and ANTs is generally $\gtrsim 0.2$ mag year-1, whereas that of CLAGNs is generally $\lesssim 0.3$ mag year-1. This may be caused by the difference between the ultraviolet light curves of TDEs/ANTs and CLAGNs, or be related to no or weak underlying AGN in TDEs/ANTs. In addition, TDEs have a redder color than ANTs at the earliest phase. Based on CVR, we selected high-probability TDE, ANT, and CLAGN candidates from MIR outbursts in samples of Jiang et al. (2021) and Masterson et al. (2024). We found that both samples are mixtures of TDEs/ANTs and CLAGNs. For MIR outbursts whose hosts are not Seyfert galaxies, we estimated that $\sim 50\%-80\%$ are TDEs and inferred a rate of infrared TDEs of $1.5-2.8\times 10^{-5}$ galaxy-1 yr-1, comparable with that of optical TDEs; The rest are CLAGNs, suggesting the presence of weak AGNs that cannot be identified using common diagnoses. Our work opens a new door for future classification of infrared selected transients based on only MIR photometric data.
@@ 1 Introduction
In quiescent galaxies, which account for the majority of galaxies, tidal disruption events (TDEs) are powerful probes of supermassive black holes (SMBHs). A TDE occurs when a star wanders too close to the SMBH for its own gravity to resist the tidal force (e.g., Rees, 1988). TDEs can be observed through bright flares lasting for months to years (e.g., van Velzen et al., 2020). Thanks to the development of time-domain surveys, $\gtrsim 100$ TDEs have been discovered. Most of TDEs are selected in the optical band (optical TDEs) via their bright UV-optical emission (e.g., van Velzen et al., 2020; Hammerstein et al., 2023), while a small number are selected in the X-ray band (X-ray TDEs) via their bright X-ray emission (e.g., Saxton et al., 2020; Sazonov et al., 2021), and some are bright in both bands (e.g., Guolo et al., 2024). According to recent sample studies, the incidence rate of optical TDEs is around $3\times 10^{-5}$ galaxy-1 yr-1 from the Zwicky Transient Facility (ZTF) sample (e.g., Yao et al., 2023), while that of X-ray TDEs is around $1\times 10^{-5}$ galaxy-1 yr-1 from the eROSITA sample (e.g., Sazonov et al., 2021; Grotova et al., 2025). In addition, the TDE rate is related to the properties of the host galaxy. For example, post-starburst galaxies have a higher incidence rate of TDEs than normal galaxies (e.g., French et al., 2016; Yao et al., 2023). Thus, the above-mentioned TDE rates may be the average value of different types of galaxies.
@@
Despite these findings, there is a lack of complete TDE samples because dust-obscured TDEs are hard to detect by optical or X-ray surveys. Fortunately, the TDE’s UV emission heats the surrounding dust to produce infrared (IR) re-radiation, known as IR echo (e.g., Jiang et al., 2016; van Velzen et al., 2016). Dust-obscured TDEs can reveal themselves by their IR echoes as IR outbursts, for example, Arp 299-B AT1 (Mattila et al., 2018) and AT 2017gbl (Kool et al., 2020). Thanks to the all-sky mid-infrared (MIR) sky survey of the Wide-field Infrared Survey Explorer (WISE, Wright et al., 2010), hundreds of MIR outbursts in centers of galaxies have been discovered (Jiang et al., 2021b; Reynolds et al., 2022; Masterson et al., 2024; Necker et al., 2025). More than $\sim$80% of these MIR outbursts were not detected in optical/X-ray bands. On the other hand, optical TDEs rarely have strong IR emission (Jiang et al., 2021a). These suggest that optical/X-ray TDEs and MIR outbursts may be two distinct groups with little overlap. The event rate of MIR outbursts obtained by Jiang et al. (2021b) is $\sim 5\times 10^{-5}$ galaxy-1 yr-1, while that by Masterson et al. (2024) and Necker et al. (2025) is lower with $\sim 2-3\times 10^{-5}$ galaxy-1 yr-1, possibly because the former work did not remove the contribution of AGN outbursts. These event rates are similar to the rate of optical and X-ray TDEs, implying that optical and X-ray surveys may have missed a significant fraction of TDEs.
@@
However, a key question remains unanswered: are these MIR outbursts TDEs? Although the authors of previously mentioned works preferred that their samples consist mainly of obscured TDEs, there is doubt (e.g., Dodd et al., 2023) that changing-look active galactic nuclei (CLAGN) contribute to these MIR outbursts significantly. CLAGNs (see review in Ricci & Trakhtenbrot, 2023) can also produce MIR outbursts (e.g., Sheng et al., 2017; Lyu et al., 2022). Despite many differences between TDEs and CLAGNs in UV/optical bands in terms of light curves (LCs) and optical spectra (e.g., Zabludoff et al., 2021), distinguishing their IR echoes remains challenging. Differences in the UV/optical LCs are blurred by the light travel time effect of the echo, and the features that may distinguishing the two are lost when the emission is reprocessed to the IR band. To make matters worse, there is a class of outbursts in active galactic nuclei (AGN) that share the characteristics of both TDEs and CLAGNs, for example AT 2016ezh (=PS16dtm, Blanchard et al., 2017; Petrushevska et al., 2023), AT 2017bgt (Trakhtenbrot et al., 2019), AT 2018bcb (=ASASSN-18jd, Neustadt et al., 2020), and some others. These outbursts are referred to as ambiguous nuclear transients (ANTs) as their natures are difficult to determine through UV/optical information: they may be TDEs that occur in AGNs (e.g., Blanchard et al., 2017), or enhanced accretions onto SMBHs (e.g., Trakhtenbrot et al., 2019). Recent MIR spectroscopy of TDEs shows that TDEs generally have strong silicate emission features, different from AGNs, which show absorption or weak emission features (Masterson et al., 2025). This implies that the silicate feature has the potential to serve as a probe to distinguish TDEs from CLAGNs. In addition, the nature of MIR outbursts can be distinguished by radio jet (e.g. for Arp 299-B AT1, Mattila et al., 2018), broad emission lines in the near-infrared (NIR) spectra (e.g. for AT 2017gbl, Kool et al., 2020) or optical spectra (e.g., Wang et al., 2022a), or shape of X-ray spectrum (e.g., Masterson et al., 2024). However, these data are not available for the vast majority of MIR outbursts. Therefore, there needs to be an effective probe to distinguish between TDEs and CLAGNs based on only MIR photometric data.
@@
Jiang et al. (2017) noticed that the dust temperature in AT 2016ezh was high with a blackbody temperature $T_{\rm BB}\gtrsim 2,000$ K initially, and then rapidly dropped to $T_{\rm BB}\sim 1,000$ K in 100–300 days. Although the temperature values here may not be accurate because the dust properties are uncertain, the trend of temperature decline is robust. Similar rapid drops in dust temperatures have also been observed in other TDEs, ANTs and TDE candidates (e.g., Mattila et al., 2018; Sun et al., 2020; Kool et al., 2020; Reynolds et al., 2022; Hinkle, 2024; Necker et al., 2025). However, such drops were not found in CLAGNs [^1] Although some works claimed a “redder when brighter” trend in MIR emission of CLAGNs (Yang et al., 2018), they did not subtract the host galaxy’s contribution to IR emission when they calculated the color. We guessed that there may be a general difference in the time variation of dust temperature between TDEs and CLAGNs. The primary purpose of this work is to verify this difference and attempt to distinguish TDEs from CLAGNs using this difference.
@@ 2 Optically-selected outbursts and data reduction
In order to study the variation of MIR colors of TDEs, CLAGNs and ANTs, and to check the disparity between different types, we collected outbursts whose type can be defined by optical observations. Although in the literature, the term ANT was also used by some authors to refer to different meanings, we will strictly use the term to mean the transient source occurring in an AGN whose nature is being debated. We sifted the samples to ensure that MIR colors can be measured with small errors. The selection criteria and the final samples are described in detail in sections 2.1 to 2.3, and the MIR properties of the sample are displayed in section 2.4.
@@ 2.1 MIR-bright TDEs, CLAGNs and ANTs
Lyu et al. (2022) has investigated the MIR properties of 68 CLAGNs reported in the literature. We started from Lyu’s sample but removed AT 2018dyk [^3] Referred to as SDSS J153308.02+443208.4 in Lyu et al. (2022) and SDSS J1115+0544. We generated the MIR LCs of the remaining 66 objects and selected bright objects with large amplitude variability in the same way as we described previously. Two CLAGNs, NGC 4151 and NGC 7582, are too bright, and their photometries are severely affected by saturated pixels, so we removed them from the sample. After setting the same criteria as for TDEs, we selected 28 MIR bright CLAGNs with significant MIR variability. As pointed out by Sheng et al. (2017), the CLAGNs in the sample have large amplitudes of MIR variability, and hence are likely to originate from changes in the accretion rate, rather than changes in obscuration or other causes.
We collected 26 ANTs [^4] We removed AT 2017gge and AT 2018dyk in Hinkle (2024) as they were classified as TDEs by us. from literatures (Frederick et al., 2021; Hinkle, 2024; Wiseman et al., 2025). We selected the sample using the same criteria and eight ANTs passed, including AT 2016ezh, AT 2017bgt, AT 2018bcb, AT 2019aalc, AT 2019avd, AT 2020agdm, AT 2021loi and AT 2022fpx.
@@ 2.2 Quiescent and outburst states
The above analysis is based on LCs generated by NEOWISE single-exposure photometries. This method of generating LCs was proved to be reliable (e.g., Jiang et al., 2021b); however, it leads to large magnitude errors for faint MIR sources. Thus, for objects with a quiescent W2 magnitude >>11, we generated LCs using time-resolved coadded data provided by the unWISE project (Meisner et al., 2018) with an image subtraction method, which was also proved reliable (e.g., Masterson et al., 2024). For each object in each band, we made a reference image using the median stack of the quiescent-state images, and then subtracted it from all the images to obtain difference images. We measured flux and error on each difference image using a PSF model whose center is fixed at the optical position for TDEs and ANTs or at the galaxy center for CLAGNs. In addition to routine statistical flux error, we considered the systematic error due to image subtraction, which is 2% of the quiescent PSF flux according to our tests using galaxies with no outbursts. We adopted the fluxes measured on difference images as excess fluxes $f_{e}$, and selected the outburst state in the same way as described previously. In Figure 2, we show an example that the data quality improves by using unWISE data. The flux errors are smaller, and more data points are selected for the outburst state, so the variation of MIR color can be better studied. However, because we have not obtained unWISE data in 2023 and beyond, we still use NEOWISE data for transients that started late, for example, AT 2022fpx, AT 2022upj, and AT 2023ugy.
@@ 2.3 Rising phase and further selections
WISEA J1003+3525 0.1189 CLAGN 26 $0.06\pm 0.06$ $0.52\pm 0.10$
WISEA J1545+1709 0.0483 CLAGN 26 $0.05\pm 0.10$ $0.71\pm 0.15$
@@ 2.4 MIR properties of the samples
In this subsection, we show the MIR properties of the optically selected transients, including $m_{\rm min,W2}$, $\delta{\rm W2}$, the peak W2 luminosity $L_{\rm max,W2}$ after quiescent level subtracted, and the rise time $t_{\rm rise}$. We calculated $L_{\rm max,W2}$ using luminosity distance $D_{L}$ converted from redshift $z$ if $z>0.01$, otherwise we used the median value of $D_{L}$ given by the NASA/IPAC Extragalactic Database [^5] https://ned.ipac.caltech.edu/. We set the start time of the outburst as between the first data point in the rising phase and the previous data point, and calculated the $t_{\rm rise}$ as the time difference between the start and the peak time in the W2 band. If the rising phase starts at the first data point of NEOWISE-R, or the outburst peaks at the last data point of the LC, we set the $t_{\rm rise}$ calculated above as the lower limit. In addition, as can be seen from Figure A2, TDE AT 2017gge and ANT AT 2018bcb peaks at the first data point of the outburst state, and hence, we could only calculate the upper limit of $t_{\rm rise}$ for them.
We show the distribution of $m_{\rm min,W2}$ and $\delta{\rm W2}$ in Figure 3(a). CLAGNs generally cause smaller amplitude of variability with $\delta{\rm W2}<1.7$ magnitude, while TDEs and ANTs can cause variation with $\delta{\rm W2}$ up to $\sim 3$ magnitude. We also show the distribution of $L_{\rm max,W2}$ and $t_{\rm rise}$ in Figure 3(b). The MIR luminosities $L_{\rm max,W2}$ of the sample are mainly in the range of $\sim 10^{42}-10^{44}$ erg s-1. The TDEs generally have lower MIR luminosities than ANTs as the median values of $L_{\rm max,W2}$ are $\sim 10^{43}$ and $\sim 10^{43.6}$ erg s-1, respectively. The rise time of the sample is distributed over a wide range between <200<200 days and >2000>2000 days. TDEs and ANTs generally have shorter rise time than CLAGNs, as among the 16 objects whose $t_{\rm rise}$ is >2>2 years, only one is TDE, and one is ANT, and the rest are CLAGNs.
Figure 3: MIR properties of the final sample (dots) and the excluded CLAGNs (crosses). The upper panel shows $m_{\rm min,W2}$ and $\delta{\rm W2}$, and the lower panel shows $L_{\rm max,W2}$ and $t_{\rm rise}$. TDEs, ANTs and CLAGNs are shown in red, green and blue symbols. The two fast-rising objects (AT 2017gge and AT 2018bcb) are marked with inverted triangles. Note that one excluded CLAGN, NGC 4395, has a luminosity of $L_{\rm max,W2}=10^{39.4}$ erg s-1 and is outside the scope of the lower panel.
@@ 3 Analysis of MIR color variation and results
This section studied the MIR color variation of the optically selected outbursts in the rising phase. We did not calculate the dust temperature as some previous works had done (e.g., Jiang et al., 2017; Reynolds et al., 2022). This is because the dust temperature depends heavily on the spectral energy distribution (SED), and its relationship with the directly observed quantity is complicated, making its error difficult to analyze. Instead, we used K-corrected MIR color W1W2W1-W2, i.e. color in the rest frame, as a probe of dust temperature. We describe how we applied the K-correction and estimated the systematic error introduced by the K-correction in Appendix B.